x+y^(1/2)=7 x^(1/2)+y=11

x+y^(1/2)=7


x^(1/2)+y=11

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

x + \sqrt{y} = 7
\sqrt{x} + y = 11
y = (7-x)^{2}
\sqrt{x} = 11 - y
x = 121 + y^{2} - 22y
x = 121 + (7-x)^{4} - 22(7-x)^{2}
x = 121 + (x^{2}-14x+49)^{2} - 22(x^{2}-14x+49)
x = 121 + (x^{4}+196x^{2}+49^{2}-28x^{3}+98x^{2}-1372x) - 22(x^{2}-14x+49)
x^{4} -28x^3 + 272x^{2}-1065x + 49^{2}-957 = 0
x^{4} -28x^3 + 272x^{2}-1065x + 1444 = 0
\Rightarrow x = 4, 3.415, 9.805, 10.78